eliminate the arbitrary constant using any method

14.y\:=\:−2C_1e^{−3x}\:+\:C_2e^{4x}


1
Expert's answer
2021-09-06T16:41:31-0400

Solution

If y(x) = −2C1e−3x + C2e4x  <=eq.1

then

y’(x) = 6C1e−3x + 4C2e4x  <= eq.2

y’’(x) = -18C1e−3x + 16C2e4x  <= eq.3

From 3* eq.1+ eq.2 => y’(x)+3 y(x) = 7 C2e4x     <= eq.4

From 3* eq.2+ eq.2 => 3 y’(x)+y’’(x) = 28 C2e4x     <= eq.5

From 4* eq.4-eq.5 => 4y’(x)+12y(x) - 3y’(x) - y’’(x) = 0  =>  y’’(x) – y’(x) – 12y(x) = 0

So the answer is y’’(x) – y’(x) – 12y(x) = 0